# Feedback-driven recurrent quantum neural network universality

### v1

@@ Abstract
Quantum reservoir computing uses the dynamics of quantum systems to process temporal data, making it particularly well-suited for learning with noisy intermediate-scale quantum devices. Early experimental proposals, such as the restarting and rewinding protocols, relied on repeating previous steps of the quantum map to avoid backaction. However, this approach compromises real-time processing and increases computational overhead. Recent developments have introduced alternative protocols that address these limitations. These include online, mid-circuit measurement, and feedback techniques, which enable real-time computation while preserving the input history. Among these, the feedback protocol stands out for its ability to process temporal information with comparatively fewer components. Despite this potential advantage, the theoretical foundations of feedback-based quantum reservoir computing remain underdeveloped, particularly with regard to the universality and the approximation capabilities of this approach. This paper addresses this issue by presenting a recurrent quantum neural network architecture that extends a class of existing feedforward models to a dynamic, feedback-driven reservoir setting. We provide theoretical guarantees for variational recurrent quantum neural networks, including approximation bounds and universality results. Notably, our analysis demonstrates that the model is universal with linear readouts, making it both powerful and experimentally accessible. These results pave the way for practical and theoretically grounded quantum reservoir computing with real-time processing capabilities.
Key Words: Recurrent Quantum Neural Network, Quantum Reservoir Computing, universality, approximation bounds, Barron system, state-space system
11footnotetext: Universität Sankt Gallen. School of Computer Science. Sankt Gallen. Switzerland. lukas.gonon@unisg.ch 22footnotetext: Imperial College. Department of Mathematics. London. United Kingdom. l.gonon@imperial.ac.uk33footnotetext: Donostia International Physics Center, Paseo Manuel de Lardizabal 4, E-20018 San Sebastián, Spain. rodrigo.martinez@dipc.org 44footnotetext: Nanyang Technological University. Division of Mathematical Sciences. School of Physical and Mathematical Sciences. Singapore. Juan-Pablo.Ortega@ntu.edu.sg
@@ 1 Introduction
Quantum reservoir computing (QRC) is a promising approach for exploiting noisy intermediate-scale quantum (NISQ) technologies for time series prediction and learning. These include ion traps, nuclear magnetic resonance, cold atoms, photonic platforms, and superconducting qubits [MMPN+21]. When implementing QRC models experimentally, it is necessary to consider the backaction and statistical effects introduced by quantum measurements. Backaction refers to the modification of a quantum state after monitoring, also known as wavefunction collapse. Due to the probabilistic nature of quantum theory, measurements must be repeated to compute the expected values of observables, which introduces a statistical component in all these methodologies. Most available experimental implementations rely on the quantum computer paradigm [DHB22, MCLCL23, SGP+22, YSK+23, CNY20, KSK+23, MDP23, PHS22, ATM25, HKB+24]. However, there is an increasing interest in extending this technique to new settings, such as optical pulses [GBGSZ23, PHGB+25], Rydberg atoms [BNGY22, KHZ+24], and quantum memristors [SMP+22, SAS+25].
Early QRC model implementations relied on the simplest possible approach, namely, the restarting protocol [DHB22, SGP+22, KSK+23, CNY20, MDP23]. In this approach, the expected values of observables are obtained by rerunning the algorithm from the first time step at each subsequent time step. This avoids the backaction effect of quantum measurements. However, the complexity of this protocol scales quadratically with the length of the input sequence, making it very time-consuming. A faster alternative is the rewinding protocol [MMPN+21, ČDLJ24], where the fading memory of the quantum reservoir is exploited to restart the algorithm with a fixed window of past time steps. This reduces the complexity of the algorithm to linear in terms of input length. Originally proposed in [CNY20], this protocol has thus far only been considered numerically [MMPG+23, ČDLJ24]. Both the restarting and rewinding protocols use repetition of previous time steps to reproduce the dynamics of the theoretical model and avoid the disruptive effect of projective measurements used to extract output information. This comes at the cost of halting the quantum dynamics at each time step and the need to buffer the input sequence. Consequently, these approaches lack one of the most important features of traditional reservoir computing, namely, the ability to process information in real time.
New protocols have been proposed to circumvent this problem. The online protocol [MMPG+23, FPL+24] uses weak measurements to find a balance between erasing and extracting information. Mid-circuit measurements and reset operations [HKB+24] can split the reservoir into two parts: memory and readout. The memory retains previous inputs, while measurements only affect the readout part. The feedback protocol [KFY24], which can be traced back to hybrid QRC techniques [PHS22, PHS23], reinjects the measured observables at each time step as parameters of an input quantum channel. This ensures that no backaction effects are present and that past input information is preserved. Note that in order to compute the observables in real time, these protocols all require several copies of the system to be run in parallel. Furthermore, these protocols can be combined with each other. For instance, the feedback protocol has been combined with both the online protocol [MSH25] and with mid-circuit measurements and reset operations [MKTG25].
Of all these approaches, the feedback protocol presents some particularly interesting features. First, the feedback protocol enables us to compute the expected values of observables from a single copy of the system by repeating one time step only. If only a few copies of the system are available, this reduces the experimental time overhead for real-time applications compared to other approaches. Second, in contrast to previous QRC models, where an erasure mechanism is added to provide fundamental properties such as the echo state property, simple unitary operations can provide these properties [KFY24]. Finally, the dynamical equations of quantum reservoirs under the feedback protocol go beyond the usual SAS paradigm of QRC models [MPO23]. These properties make the feedback protocol a promising candidate for exploring QRC applications.
@@
Despite all these interesting features of QRC models based on the feedback protocol, basic theoretical results such as approximation bounds, universality proofs, and criteria for these architectures to satisfy the echo state or the fading memory properties do not exist yet. Moreover, previous universality results concerning QRC models have relied on the use of polynomial output layers [CN19, CNY20, NMPG+21, SMPS+24, SGLZ24], which yield a polynomial algebra that can then be used with the Stone-Weierstrass theorem to obtain universality statements. Nevertheless, most numerical and experimental implementations of reservoir computers use linear output layers due to their simplicity and fast training.
In this paper, we address these theoretical shortcomings by proposing and analyzing a recurrent quantum neural network (RQNN) architecture that builds on and extends the feedforward quantum neural networks (QNN) introduced in [GJ25] to a dynamic quantum reservoir computing setting. In the context of the feedback protocol, we derive approximation error bounds and prove universality statements for RQNN families with a linear output layer. Universality refers to the ability of these families to uniformly approximate arbitrarily well a large category of dynamic processes, namely fading memory input / output systems (defined later on). It is important to note that, strictly speaking, these findings apply exclusively to variational quantum circuits for which all parameters are trained and are only tentative for quantum reservoir systems in which some parameters in the recurrent layer are randomly generated. More explicitly, in the variational method, the parameters of the quantum circuit are tuned alongside the final readout layer unlike the reservoir method which usually involves randomly assigning circuit parameters and optimising only the output layer, making it a simpler technique. Most previous literature on RC and QRC universality [GO18b, GO18a, GO20, GO21, CN19, CNY20, NMPG+21, SMPS+24, SGLZ24] implicitly assumes the search for an optimal model within a class in which all parameters are estimated. This is different from the approach taken in, for example, [GGO23], where it is explicitly assumed that the circuit parameters are randomly sampled. Finally, the error bounds and the universality results presented in this paper deal with the approximation properties of RQNNs but do not explore their behavior in relation to estimation errors. That will be the subject of a future work. The reader interested in this question is encouraged to check with [GGO20, CAM25].
@@
The structure of the paper is as follows. Section 2.1 describes the RQNN model, which is an extension of the static QNN introduced in [GJ25] with state feedback. Section 2.2 generalizes the approximation bounds obtained in [GJ25] to accommodate first derivatives and uses these results (see Proposition 2.5 and Corollary 2.6) to study the properties of the RQNN state maps in the uniform approximation of more general state equations as well as in a square-integrable sense. These results are then used in Section 2.3 to prove the universal uniform approximation properties of the filters associated with RQNN systems. More specifically, in Proposition 2.7 we provide filter approximation bounds that show that RQNNs are able to uniformly approximate the filters induced by any contracting Barron-type state-space system. Finally, Theorem 2.9 of Section 2.4 extends this universality property to the much larger category of arbitrary fading memory, causal, and time-invariant filters. These results show that RQNNs have approximation properties as competitive as those of popular reservoir computing/state-space system families like echo state networks [GO18a, GO20, GO21, GGO23], state-affine systems [GO18b, GO20], or linear systems with polynomial / neural network readouts [GO18b]. The paper concludes with Section 3, where the main contributions and outlook of the paper are summarized.
@@
Reservoir computing definitions. Let the symbol $(\mathbb{R}^{n})^{\mathbb{Z}}$ denote the set of infinite real sequences of the form $\underline{\bm{z}}=(\ldots,\bm{z}_{-1},\bm{z}_{0},\bm{z}_{1},\ldots)$, $\bm{z}_{i}\in\mathbb{R}^{n}$, $i\in\mathbb{Z}$; $(\mathbb{R}^{n})^{\mathbb{Z}_{-}}$ is the subspace consisting of left infinite sequences: $(\mathbb{R}^{n})^{\mathbb{Z}_{-}}=\{\underline{\bm{z}}=(\ldots,\bm{z}_{-2},\bm{z}_{-1},\bm{z}_{0})\mid\bm{z}_{i}\in\mathbb{R}^{n},i\in\mathbb{Z}_{-}\}$. Analogously, $(D_{n})^{\mathbb{Z}}$ and $(D_{n})^{\mathbb{Z}_{-}}$ stand for infinite and semi-infinite sequences, with elements in the subset $D_{n}\subset\mathbb{R}^{n}$. Let $D_{n}\subset\mathbb{R}^{n}$ and $B_{N}\subset\mathbb{R}^{N}$. We refer to the maps of the type $U:(D_{n})^{\mathbb{Z}}\longrightarrow(B_{N})^{\mathbb{Z}}$ as filters and to those like $H:(D_{n})^{\mathbb{Z}}\longrightarrow B_{N}$ (or $H:(D_{n})^{\mathbb{Z}_{-}}\longrightarrow B_{N}$) as functionals. A filter $U:(D_{n})^{\mathbb{Z}}\longrightarrow(B_{N})^{\mathbb{Z}}$ is called causal when for any two elements $\bm{z},\bm{w}\in(D_{n})^{\mathbb{Z}}$ that satisfy that $\bm{z}_{\tau}=\bm{w}_{\tau}$ for any $\tau\leq t$, for a given $t\in\mathbb{Z}$, we have that $U(\bm{z})_{t}=U(\bm{w})_{t}$. Let $T_{\tau}:(D_{n})^{\mathbb{Z}}\longrightarrow(D_{n})^{\mathbb{Z}}$, $\tau\in\mathbb{Z}$ be the time delay operator defined by $T_{\tau}(\bm{z})_{t}:=\bm{z}_{t-\tau}$. The filter $U$ is called time-invariant when it commutes with the time delay operator, that is, $T_{\tau}\circ U=U\circ T_{\tau}$, for any $\tau\in\mathbb{Z}$, where the two operators $T_{\tau}$ are defined in the appropriate sequence spaces. Finally, we recall there is a bijection between causal time-invariant filters and functionals on $(D_{n})^{\mathbb{Z}_{-}}$, and we can use them interchangeably [GO18a].
@@
Consider a recurrent neural network determined by two maps, namely the recurrent layer or the state map $F:\mathbb{R}^{N}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}^{N}$, $n,N\in\mathbb{N}$, and a readout or observation map $h:\mathbb{R}^{N}\rightarrow\mathbb{R}^{m}$, $m\in\mathbb{N}$, given by
@@
where $t\in\mathbb{Z}$, $\bm{z}_{t}$ denotes the input, $\bm{x}_{t}\in\mathbb{R}^{N}$ is the state vector, and ${\bf y}_{t}\in\mathbb{R}^{m}$ is the output vector. If $F$ is a fixed map that has no parameters that are estimated during the training process, the recurrent layer is referred to as a reservoir and the recurrent neural network system (1) as a reservoir computing system.
@@
Consider now subsets $B_{N}\subset\mathbb{R}^{N}$ and $D_{n}\subset\mathbb{R}^{n}$ and a recurrent layer defined on them, that is, $F:B_{N}\times D_{n}\longrightarrow B_{N}$ and $h:B_{N}\rightarrow\mathbb{R}^{m}$. Denote by $D_{m}:=h(B_{N})\subset{\mathbb{R}}^{m}$. The recurrent system $F$ is said to have the echo state property with respect to inputs in $\left(D_{n}\right)^{\mathbb{Z}}$ when for any $\underline{\bm{z}}\in\left(D_{n}\right)^{\mathbb{Z}}$ there exists a unique element $\underline{\bm{x}}\in\left(B_{N}\right)^{\mathbb{Z}}$ that satisfies the first equation in (1), for each $t\in\mathbb{Z}$. When the echo state property holds, a unique filter $U^{F}:(D_{n})^{\mathbb{Z}}\longrightarrow(B_{N})^{\mathbb{Z}}$ can be associated to the recurrent system determined by $F$, namely, $U^{F}(\bm{z})_{t}:=\bm{x}_{t}\in B_{N}$, for all $t\in\mathbb{Z}$. We will denote by $U^{F}_{h}:(D_{n})^{\mathbb{Z}}\longrightarrow(D_{m})^{\mathbb{Z}}$ the corresponding filter determined by the entire recurrent system, that is, $U^{F}_{h}(\bm{z})_{t}=h\left(U^{F}(\bm{z})_{t}\right):={\bf y}_{t}\in D_{m}$, for all $t\in\mathbb{Z}$. The filters $U^{F}$ and $U^{F}_{h}$ are causal and time-invariant by construction. The echo state property is much related with the so-called fading memory property defined as the continuity of $U^{F}_{h}$ with respect to weighted norms in its domain and codomain [BC85] or the product topologies when $D_{n}$ and $D_{m}$ are compact [GO18a]. It can be shown that when $D_{m}$ is compact, the echo state property implies the fading memory property [Man20, OR25b]; see [OR25a] for a comprehensive account of the dynamical implications of the fading memory property.
2 Recurrent quantum neural network universality
This section contains approximation guarantees and universality results for the recurrent quantum neural network (RQNN) family. To achieve this, we will first introduce in detail in Section 2.1 the recurrent quantum neural networks employed. Then, in Section 2.2 we prove refined approximation error bounds (that generalize those in [GJ25]) for feedforward quantum neural networks (QNNs) that allow us to control the error committed when approximating a function and its derivatives simultaneously. These error bounds show how recurrent QNNs can be used to approximate state-space maps arbitrarily well as long as these are sufficiently smooth and satisfy Barron-type conditions; RQNN state maps are hence universal in that category. These bounds are devised with respect to $L^{\infty}$ and $L^{2}$-type norms. As we shall prove, in the $L^{\infty}$ case, the universality of the RQNN family still holds with respect to state maps that do not necessarily satisfy the Barron condition, even though in that framework we do not formulate approximation bounds. Finally, in the last two sections, we exploit all these results on the approximation of state maps to obtain universality statements and error bounds for the approximation of arbitrary causal, time-invariant, and fading memory filters using a modified recurrent QNN. In addition to the tools developed here, our proofs of these results rely on techniques from [GO20, GO21] and the overall strategy is reminiscent of the so-called internal approximation approach introduced in [GO18a, Theorem 3.1 (iii)] for echo state networks, which consists of obtaining approximation results for filters out of statements of that type for the state maps that generate them.
2.1 The recurrent quantum neural network (RQNN) construction

### v2

@@ Abstract
Quantum reservoir computing uses the dynamics of quantum systems to process temporal data, making it particularly well-suited for machine learning with noisy intermediate-scale quantum devices. Recent developments have introduced feedback-based quantum reservoir systems, which process temporal information with comparatively fewer components and enable real-time computation while preserving the input history. Motivated by their promising empirical performance, in this work, we study the approximation capabilities of feedback-based quantum reservoir computing. More specifically, we are concerned with recurrent quantum neural networks, which are quantum analogues of classical recurrent neural networks. Our results show that regular state-space systems can be approximated using quantum recurrent neural networks without the curse of dimensionality and with the number of qubits only growing logarithmically in the reciprocal of the prescribed approximation accuracy. Notably, our analysis demonstrates that quantum recurrent neural networks are universal with linear readouts, making them both powerful and experimentally accessible. These results pave the way for practical and theoretically grounded quantum reservoir computing with real-time processing capabilities.
@@ 1 Introduction
Recent advances in quantum computing have led to a rapid development of quantum machine learning methods. These methods aim to exploit the potential computational speed-up and reduced complexity offered by quantum computing for machine learning purposes. For learning problems with temporal structure, quantum reservoir computing (QRC) has emerged as a promising approach for exploiting noisy intermediate-scale quantum (NISQ) technologies. In contrast to classical machine learning methods based on bits valued in $\{0,1\}$, quantum bits (qubits) can be in a continuum of states. QRC aims to exploit this fundamental difference to build efficient machine learning methods for time series prediction and learning.
In this paper, we are concerned with recurrent quantum neural networks (RQNN), a particular type of quantum reservoir computing method. RQNNs are a quantum analogue to classical recurrent neural networks. RQNNs are built from quantum neural networks (QNNs), with weights and biases typically realized via quantum circuits. Thus, these networks can be evaluated directly on quantum computers. Thereby, quantum machine learning aims to achieve a significant increase in neural network expressivity and computational speed-up in inference and training.
Motivated by their promising empirical performance, in this work, we study the approximation capabilities of feedback-based quantum reservoir computing methods and, specifically, RQNNs. In particular, our work provides precise bounds on the number of qubits and the size of the underlying quantum circuit that is required to guarantee a prescribed approximation accuracy. Our results show that QRNNs can approximate regular state-space systems using a quantum circuit with qubit number only growing logarithmically in the reciprocal of the prescribed approximation accuracy and with error rates not suffering from the curse of dimensionality. Thereby, our results pave the way for theoretically grounded quantum reservoir computing with real-time processing capabilities.
1.1 Related Literature
Quantum reservoir computing methods have been extensively studied for a variety of time-series prediction and learning tasks, employing different architecture types such as online protocols (Mujal et al., 2023; Franceschetto et al., 2024), mid-circuit measurements and reset operations (Hu et al., 2024; Murauer et al., 2025), feedback protocols (Kobayashi et al., 2024), QRC with quantum memristors (Spagnolo et al., 2022) and hybrid QRC techniques (Pfeffer et al., 2022; 2023). We provide a detailed discussion of QRC methods in Appendix A.
Despite these promising developments, key questions regarding universal approximation capabilities and expressivity of feedback-driven QRC methods have not been addressed in the literature. For classical neural networks, qualitative and quantitative universal approximation theorems have been extensively studied, with seminal works including, e.g.  Hornik (1991); Barron (1993); Yarotsky (2017). Universality results for the dynamic reservoir computing setting have been obtained in (Grigoryeva & Ortega, 2018a; b; Gonon & Ortega, 2020; 2021; Gonon et al., 2023) for echo state networks, state-affine systems and linear systems with polynomial / neural network readouts. For (feedforward) QNNs first qualitative results on universal approximation properties of QNNs have been proved only very recently Pérez-Salinas et al. (2020); Schuld et al. (2021). Subsequently, quantitative approximation error bounds for feedforward QNNs were proved in Gonon & Jacquier (2025); Yu et al. (2024); Aftab & Yang (2024).
@@ 1.1 Related Literature
For RQNNs, no quantitative approximation error bounds have been previously available in the literature. Moreover, previous universality results concerning QRC models have relied on the use of polynomial output layers (Chen & Nurdin, 2019; Chen et al., 2020; Nokkala et al., 2021; Sannia et al., 2024b; a), which yield a polynomial algebra that can then be used with the Stone-Weierstrass theorem to obtain universality statements. Nevertheless, most numerical and experimental implementations of reservoir computers use linear output layers due to their simplicity and fast training.
1.2 Contributions
For applications of QRC methods in learning tasks with temporal dependence, a precise understanding of RQNN approximation capabilities is essential. In this paper, we derive approximation error bounds and prove universality statements for RQNN families with a linear output layer and in the context of the feedback protocol. Universality refers to the ability of these families to uniformly approximate arbitrarily well a large category of dynamic processes, so-called fading memory input/output systems. Thereby, we contribute to a precise understanding of RQNN approximation capabilities in several aspects.
  • We provide RQNN approximation error bounds for regular state-space systems. Our first main result, Theorem 4.6, shows that RQNNs are able to approximate regular state-space systems without the curse of dimensionality, using quantum circuits with qubit number only growing logarithmically in the reciprocal of the prescribed approximation accuracy.
  • In our second main result, Theorem 4.8, we prove that RQNNs can uniformly approximate the arbitrary fading memory, causal, and time-invariant filters. In particular, RQNNs have approximation properties as competitive as those of popular reservoir computing/state-space system families like echo state networks, state-affine systems, or linear systems with polynomial/neural network readouts.
  • To prove these results, we first derive novel qualitative and quantitative approximation error results for using feedforward QNNs to approximate functions and their derivatives (see Proposition 4.4 and Corollary 4.5).
In comparison to Gonon & Jacquier (2025), our RQNNs introduce memory through a feedback loop. Mathematically analysing our RQNNs architecture hence requires a novel, intricate analysis of QNN approximations of functions jointly with their derivatives. Moreover, approximation analysis in the temporal domain is inherently much more challenging due to the feedback loop. Proving Theorems 4.6 and 4.8 thus requires new techniques specifically tailored to deal with this situation (see Appendix C). Most previous literature on RC and QRC universality (Grigoryeva & Ortega, 2018a; b; Gonon & Ortega, 2020; 2021; Chen & Nurdin, 2019; Chen et al., 2020; Nokkala et al., 2021; Sannia et al., 2024b; a) implicitly assumes the search for an optimal model within a class in which all parameters are estimated. Also our results are formulated for variational quantum circuits for which all parameters are trainable. Nevertheless, the obtained results and developed proof techniques also promise to be useful for QRC systems in which certain parameters in the recurrent layer are randomly generated. Our RQNN architecture builds on and extends the feedforward QNN architecture introduced in Gonon & Jacquier (2025), which also admits results for the randomized setting. Hence, combining the techniques developed here with these randomized architectures may provide fruitful for studying randomization in the dynamic quantum reservoir computing setting. Moreover, the obtained approximation error bounds may serve as a crucial ingredient for bounding the overall generalization error of QRC methods, by combining our results with suitable risk bounds as obtained in other contexts in Gonon et al. (2020); Chmielewski et al. (2025).
1.3 Outline
@@
The paper is structured as follows. Section 2 introduces background on filters, functionals, fading-memory and echo state properties. Section 3 describes the RQNN model, a recurrent QNN with state feedback. Section 4.1 derives QNN approximation error bounds for functions and their first derivatives. We then use these results (see Proposition 4.4 and Corollary 4.5) to study the properties of the RQNN state maps for approximating more general state equations. These results are then used in Section 4.2 to prove the universal uniform approximation properties of the filters associated with RQNN systems. More specifically, in Theorem 4.6 we provide filter approximation bounds that show that RQNNs can uniformly approximate the filters induced by any contracting Barron-type state-space system. Finally, Theorem 4.8 of Section 4.3 extends this universality property to the much larger category of arbitrary fading memory, causal, and time-invariant filters. The paper concludes with Section 5, where the main contributions and outlook of the paper are summarized.
2 Background on filters and functionals
We start by introducing the input-output maps to be learnt in the dynamic setting. In a static context, input-output maps are given by functions of the form $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$. For learning with temporal dependence, the relevant input-output maps are filters and functionals defined on sequences.
@@ 2 Background on filters and functionals
Specifically, let $(\mathbb{R}^{n})^{\mathbb{Z}}$ denote the set of infinite real sequences of the form $\underline{\bm{z}}=(\ldots,\bm{z}_{-1},\bm{z}_{0},\bm{z}_{1},\ldots)$, $\bm{z}_{i}\in\mathbb{R}^{n}$, $i\in\mathbb{Z}$; $(\mathbb{R}^{n})^{\mathbb{Z}_{-}}$ is the subspace consisting of left infinite sequences: $(\mathbb{R}^{n})^{\mathbb{Z}_{-}}=\{\underline{\bm{z}}=(\ldots,\bm{z}_{-2},\bm{z}_{-1},\bm{z}_{0})\mid\bm{z}_{i}\in\mathbb{R}^{n},i\in\mathbb{Z}_{-}\}$. Analogously, $(D_{n})^{\mathbb{Z}}$ and $(D_{n})^{\mathbb{Z}_{-}}$ stand for infinite and semi-infinite sequences, with elements in the subset $D_{n}\subset\mathbb{R}^{n}$. Let $D_{n}\subset\mathbb{R}^{n}$ and $B_{N}\subset\mathbb{R}^{N}$. We refer to the maps of the type $U:(D_{n})^{\mathbb{Z}}\longrightarrow(B_{N})^{\mathbb{Z}}$ as filters and to those like $H:(D_{n})^{\mathbb{Z}}\longrightarrow B_{N}$ (or $H:(D_{n})^{\mathbb{Z}_{-}}\longrightarrow B_{N}$) as functionals. A filter $U:(D_{n})^{\mathbb{Z}}\longrightarrow(B_{N})^{\mathbb{Z}}$ is called causal when for any two elements $\bm{z},\bm{w}\in(D_{n})^{\mathbb{Z}}$ that satisfy that $\bm{z}_{\tau}=\bm{w}_{\tau}$ for any $\tau\leq t$, for a given $t\in\mathbb{Z}$, we have that $U(\bm{z})_{t}=U(\bm{w})_{t}$. Let $T_{\tau}:(D_{n})^{\mathbb{Z}}\longrightarrow(D_{n})^{\mathbb{Z}}$, $\tau\in\mathbb{Z}$ be the time delay operator defined by $T_{\tau}(\bm{z})_{t}:=\bm{z}_{t-\tau}$. The filter $U$ is called time-invariant when it commutes with the time delay operator, that is, $T_{\tau}\circ U=U\circ T_{\tau}$, for any $\tau\in\mathbb{Z}$, with the two operators $T_{\tau}$ defined in the appropriate sequence spaces. Finally, there is a bijection between causal time-invariant filters and functionals on $(D_{n})^{\mathbb{Z}_{-}}$, and we can use them interchangeably (Grigoryeva & Ortega, 2018b).
@@
A specific class of filters is given by state-space systems (such as recurrent neural networks) determined by two maps, namely the recurrent layer or the state map $F:\mathbb{R}^{N}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}^{N}$, $n,N\in\mathbb{N}$, and a readout or observation map $h:\mathbb{R}^{N}\rightarrow\mathbb{R}^{m}$, $m\in\mathbb{N}$, given by
@@
where $t\in\mathbb{Z}$, $\bm{z}_{t}$ denotes the input, $\bm{x}_{t}\in\mathbb{R}^{N}$ is the state vector, and ${\bf y}_{t}\in\mathbb{R}^{m}$ is the output vector.
@@
Consider now subsets $B_{N}\subset\mathbb{R}^{N}$ and $D_{n}\subset\mathbb{R}^{n}$ and a recurrent layer defined on them, that is, $F:B_{N}\times D_{n}\longrightarrow B_{N}$ and $h:B_{N}\rightarrow\mathbb{R}^{m}$. Denote by $D_{m}:=h(B_{N})\subset{\mathbb{R}}^{m}$. The recurrent system $F$ is said to have the echo state property with respect to inputs in $\left(D_{n}\right)^{\mathbb{Z}}$ when for any $\underline{\bm{z}}\in\left(D_{n}\right)^{\mathbb{Z}}$ there exists a unique element $\underline{\bm{x}}\in\left(B_{N}\right)^{\mathbb{Z}}$ that satisfies the first equation in (1), for each $t\in\mathbb{Z}$. When the echo state property holds, a unique filter $U^{F}:(D_{n})^{\mathbb{Z}}\longrightarrow(B_{N})^{\mathbb{Z}}$ can be associated to the recurrent system determined by $F$, namely, $U^{F}(\bm{z})_{t}:=\bm{x}_{t}\in B_{N}$, for all $t\in\mathbb{Z}$. We will denote by $U^{F}_{h}:(D_{n})^{\mathbb{Z}}\longrightarrow(D_{m})^{\mathbb{Z}}$ the corresponding filter determined by the entire recurrent system, that is, $U^{F}_{h}(\bm{z})_{t}=h\left(U^{F}(\bm{z})_{t}\right):={\bf y}_{t}\in D_{m}$, for all $t\in\mathbb{Z}$. The filters $U^{F}$ and $U^{F}_{h}$ are causal and time-invariant by construction. The echo state property is much related with the so-called fading memory property defined as the continuity of $U^{F}_{h}$ with respect to weighted norms in its domain and codomain (Boyd & Chua, 1985) or the product topologies when $D_{n}$ and $D_{m}$ are compact (Grigoryeva & Ortega, 2018b). It can be shown that when $D_{m}$ is compact, the echo state property implies the fading memory property (Manjunath, 2020; Ortega & Rossmannek, 2025b); see Ortega & Rossmannek (2025c) for a comprehensive account of the dynamical implications of the fading memory property as well as Ortega & Rossmannek (2025a) for a stochastic version.
3 Recurrent quantum neural network architecture
[…]